OpenStudy (tiffany):

P 13 _____ = _______ 8 2

OpenStudy (amistre64):

cross multiply; really a memory device and not a real math solution... the real solution is to multiply everything by 8 8p 8(13) --- = ---- 8 2 p = 4(13) = 40+12 = 52 maybe?